# Graph Condensation: A Survey

### v1

@@ 4 Graph Condensation Methods
In this section, we dive into the specifics of GC and conduct a systematic analysis of four distinct GC methodologies: effective GC, generalized GC, fair GC, and efficient GC, exploring the techniques employed to achieve their respective objectives.
@@ 4.1 Effective Graph Condensation
As mentioned in Section 2, we further divide effective GC methods into two categories according to the enhanced components within GC framework, including optimization-driven methods and module-enhanced methods.
@@ 4.1.1 Optimization-Driven Methods
Designing more advanced optimization strategies is often considered the critical driving force to achieve effective GC, thus steering many studies in this field. The first GC method Jin et al. (2022b) employs the gradient matching as the optimization strategy. However, the limitations associated with single-step gradient alignment constrain the overall performance of GC. Therefore, the follow-up studies are proposed to explore and enhance the optimization process. For example, CTRL Anonymous (2023) argues that the distance measurement in gradient matching is coarse, which relies on cosine similarity and solely focuses on the directional aspect of gradients. This fails to adequately capture the vector characteristics of gradients, thereby introducing biases in the gradient matching process. Therefore, CTRL proposes to refine the matching objective by adding the magnitudes distance. Furthermore, CTRL empirical discovers that gradient magnitude matching not only explicitly accounts for the vector properties of gradients, but also implicitly benefits the alignment of frequency distributions between condensed and original graphs. This provides another compelling rationale for the performance improvement.
@@
Despite the improvements from fine-grained gradient matching, CTRL is still focuses on short-range matching. This leads to the short-sight issue and fails to capture the holistic GNN learning behaviors, potentially compromising the quality of condensed graphs. To address this limitation, SFGC Zheng et al. (2023c) introduces trajectory matching in GC and proposes to align the long-term GNN learning behaviors between the original graph and condensed graph. As a result, it enables a more comprehensive and efficient knowledge transfer to the condensed graph.
@@
Beyond aligning gradients or trajectories during model training, kernel ridge regression is employed for precise solutions in bi-level optimization objectives. For instance, GC-SNTK Wang et al. (2023) discards the GNN and employ the kernel ridge regression, incorporating graph neural tangent kernels Du et al. (2019) as the relay model. This modification not only facilitates the first-order optimization of the condensed graph, but also eliminates the requirement for multiple initializations of relay GNN, resulting in a more stable condensation procedure. KIDD Xu et al. (2023) employs kernel ridge regression for graph classification tasks and further improves the method by removing the non-linear activation in graph neural tangent kernels. This approach not only simplifies the matrix multiplications in the iterative updates of condensed graphs but also leads to substantial improvements in model accuracy compared to the gradient matching method.
4.1.2 Module-Enhanced Methods
@@
Optimization-driven methods enhance the effectiveness of GC by emphasizing the optimization procedure. In contrast, other studies improve the performance by augmenting the other modules with in GC framework.
As a representative method in this category, MSGC Gao and Wu (2023) improves the GC from graph structure construction. Unlike many GC approaches which synthesize a single dense adjacency matrix, MSGC instead pre-defines multiple small-scale sparse graphs. These pre-defined structures allow each condensed node encompasses distinct neighborhoods. The variation in structure enables GNNs to capture a broader information, leading to diverse node embeddings and enhanced performance. PUMA Liu et al. (2023c) emphasises the significance of label quantity in GC and improves the condensation quality by incorporating pseudo labels from the original graph. The inclusion of additional supervision signals enhances class representation and ensures more accurate information preservation in the condensed graph. Moreover, GroC Li et al. (2023) designs a perturbation module in the process of generating condensed graphs. This module maximizes the distance between the condensed graph and the original graph by selectively perturbing the condensed graph that are insufficiently informative. By utilizing adversarial training technique, these perturbations are capable of exploring the neighborhood areas within the parameter space of the condensed graph and significantly enhancing the gradient matching process.
@@ 4.2 Generalized Graph Condensation
The essence of generalization lies in preserving crucial task-related information while mitigating bias injection into the condensed graph. Nevertheless, some GNN relay models, employed for encoding both structure and attribute information, inevitably result in a loss of structural information and create an entanglement between the condensed graph and the relay model. These issues significantly affect the generalization ability to unseen models and tasks. To explore the cause, GCEM Liu et al. (2023a) conducts an investigation across various GNN architectures, including different spatial and spectral GNNs. They find that the performance gap increases when the relay model and downstream model have different architectures. This is due to that distinct spectral preferences of GNNs have varying impacts on the significance of different graph eigenvectors Bo et al. (2023); Zhu et al. (2021), consequently leading to alterations in node features. For instance, if a GNN functions primarily as a low-pass filter, it will prioritize low-frequency information in node features, overlooking the full spectrum. The loss of high-frequency information will compromise the high-pass filter GNNs’ performance.
To address this issue, generalized GC methods are developed from the spectral perspective. For example, GCEM Liu et al. (2023a) circumvents the conventional relay GNN and directly generates the eigenbasis for the condensed graph, thereby eliminating the spectrum bias inherent in relay GNNs. To address the inconsistencies in eigenvector quantity and dimension, GCEM aligns the condensed eigenvectors with the selected eigenvectors of the original graph and matches the class features in the subspace defined by these eigenvectors.
@@
Apart from leveraging the advanced relay model, SGDD Yang et al. (2023a) employs the Laplacian energy distribution (LED) to analyze the structural properties of the graph, quantifying the spectral shift between condensed and original graphs. A substantial LED shift suggests that crucial structural information from the original graph is lost, diminishing the generalization capabilities of the condensed graph. Therefore, SGDD leverages the LED as a structural shift metric and designs an additional regularization to enhance the structure of the condensed graph. This regularization aligns the LED between the original graph and the condensed graph, facilitating the broadcast of structural information. By preserving crucial structural information, SGDD effectively reduces the performance discrepancies observed across different GNN architectures and various tasks, e.g. node classification, anomaly detection and link prediction.
@@ 4.3 Fair Graph Condensation
GC can effectively capture the essential task-specific characteristics of the original graph by exclusively optimizing for task performance. However, this task-centric optimization approach may introduce fairness concerns, primarily attributed to the indiscriminate transmission of information into the condensed graph. Especially in scenarios with intensive compression rates, GC tends to amplify biases as it prioritizes the preservation of predominant and representative information. Consequently, GNNs trained on condensed graphs exhibit more severe fairness issues than those trained on original graphs Feng et al. (2023); Mao et al. (2023), which challenges the applicability of GC in high-stake applications Mehrabi et al. (2021); Feng et al. (2023).
@@
To mitigate bias in condensed graphs, FGD Feng et al. (2023) initially investigates the relationship between the feature spaces of original graph and condensed graph, and constructs an estimator for sensitive attributes. Following this, the variance of the estimated sensitive group membership is employed as the bias measurement and integrated as the regularization to facilitate the unbiased condensed graph. Moreover, GCARe Mao et al. (2023) directly constrains the condensation process with adversarial training. It additionally incorporates a discriminator into the relay model, which is designed to predict the sensitive groups of node representations. The relay model operates as a generator, generating the fair data to deceive the discriminator. Consequently, the condensation process is formulated as a min-max optimization problem, striking a balance between data generation and debiasing efforts.
4.4 Efficient Graph Condensation
The GC process requires significant computational resources. To figure out the root causes of heavy computation, we break down the GC process into three fundamental stages, including: efficient graph encoding, efficient optimization and efficient graph generation, and examine acceleration methods for each component.
4.4.1 Efficient Graph Encoding
In GC, the graph encoding process is executed by the relay model. Generally, relay models used in GC methods adopt either the message-passing framework or the graph neural tangent kernel framework. All these frameworks stack multiple propagation and transformation operations to aggregate and update information according to the graph structure. However, as multiple propagation processes are executed, the number of aggregated neighboring nodes increases exponentially, leading to significant computational demands.
To mitigate the expensive computation resulting from propagation, some GC methods borrow the idea from SGC Wu et al. (2019) and remove the non-linear transformation layers between aggregation layers in the relay model. For example, Gcond Jin et al. (2022b) and PUMA Liu et al. (2023c) utilize SGC to encode the original graph. KIDD Xu et al. (2023) simplifies the graph neural tangent kernel by removing non-linear activation at certain layers. By this means, the propagation of these methods can be pre-computed in the pre-processing stage and only needs to be executed once. Instead of performing propagation during each training epoch, the online time complexity of GC is significantly reduced than utilizing other relay models.
4.4.2 Efficient Optimization
In GC, the optimization procedure often requires numerous iterations to update the relay model before optimizing the condensed graph, leading to considerable computation.
To mitigate this problem, DosCond Jin et al. (2022a) proposes the one-step gradient matching strategy which discards the relay model update and only performs gradient matching for one single step. The theoretical analysis posits that minimizing the one-step matching loss serves a dual purpose: it not only directs the learning process of condensed graphs towards lower losses on original graphs but also delineates the trajectory for updating the condensed data. Moreover, one-step matching eliminates the need for hyper-parameter tuning, such as the number of iterations of bi-level optimization, further simplifying the condensation process.
Instead of matching gradients, distribution matching methods change the optimization strategy towards approximating original graph in the feature space. This approach leads to a shift in the role of the relay model within GC. It transitions from aligning task performances to transforming both original and condensed graphs into a unified space for distribution alignment. Consequently, distribution matching obviates the need to compute the gradients and trajectories w.r.t. parameters of the relay model in each training iteration, leading to a substantial reduction in computational complexity. Specifically, GCDM Liu et al. (2022), CaT Liu et al. (2023b), PUMA Liu et al. (2023c) and GCEM Liu et al. (2023a) compute node distributions for both original and condensed graphs in the same feature space and employ the maximum mean discrepancy Gretton et al. (2012) between each pair of class distributions as the optimization objective.
Another research line to avoid the bi-level optimization is kernel ridge regression-based methods Xu et al. (2023); Wang et al. (2023). These methods reformulate the classification task as a kernel ridge regression task and incorporate the graph neural tangent kernel within kernel ridge regression to encode the original graph. Kernel ridge regression provides an efficient closed-form solution, dispensing with the necessity for iterative computations in updating the relay model and resulting in significantly faster convergence speeds.
4.4.3 Efficient Graph Generation
Condensed graph generation targets two components: the feature matrix and the adjacency matrix. Modeling these matrices typically requires two distinct optimization processes. For example, GCond Jin et al. (2022b) models the condensed structure as a function of the condensed nodes, requiring the optimization of an additional adjacency matrix generator. To simplify the optimization procedure, several studies Jin et al. (2022b); Zheng et al. (2023c); Liu et al. (2022, 2023b, 2023c) discard the adjacency matrix and instead use an identity matrix to represent condensed structure. These structure-free GC methods implicitly encode the original graph’s topological structure into discriminative node attributes via relay GNNs and these enriched embeddings are informative to train different downstream GNNs.

### v2

@@ 4 Graph Condensation Methods
In this section, we dive into the specifics of GC and conduct a systematic analysis of five distinct GC methodologies: effective GC, generalized GC, efficient GC, fair GC, and robust GC, exploring the techniques employed to achieve their respective objectives. In addition to our proposed taxonomy, we also provide a detailed summary of GC methods, highlighting their distinct characteristics, in Table II. This table offers a comprehensive comparison across several dimensions, including category, optimization strategy, condensed graph generation method, and downstream tasks.
@@ 4.1 Effective Graph Condensation
As mentioned in Section 2, we further divide effective GC methods into two categories according to the enhanced components within the GC framework, including optimization-driven methods and augmentation-based methods.
@@ 4.1.1 Optimization-Driven Methods
Designing more advanced optimization strategies is often considered the critical driving force to achieve effective GC, thus steering many studies in this field. The first GC method, GCond [25], employs gradient matching as the optimization strategy. However, the limitations associated with single-step gradient alignment constrain the overall performance of GC. Therefore, follow-up studies are proposed to explore and enhance the optimization process.
Gradient matching. CTRL [51] argues that the distance measurement in gradient matching is coarse, which relies on cosine similarity and solely focuses on the directional aspect of gradients. This fails to adequately capture the vector characteristics of gradients, thereby introducing biases in the gradient matching process. Therefore, CTRL proposes to refine the matching objective by adding the magnitudes distance. Furthermore, CTRL empirical discovers that gradient magnitude matching not only explicitly accounts for the vector properties of gradients, but also implicitly benefits the alignment of frequency distributions between condensed and original graphs. This provides another compelling rationale for the performance improvement.
@@
Trajectory matching. Despite the improvements from fine-grained gradient matching, CTRL still focuses on short-range matching. This leads to the short-sight issue and fails to capture the holistic GNN learning behaviors, potentially compromising the quality of condensed graphs. To address this limitation, SFGC [28] introduces trajectory matching in GC and proposes to align the long-term GNN learning behaviors between the original graph and condensed graph. As a result, SFGC enables a more comprehensive knowledge transfer to the condensed graph. Furthermore, GEOM [59] identifies the restricted supervision signals in trajectory matching and argues that difficult nodes are key to the performance gap of GNNs trained on the condensed graph. To address this, GEOM evaluates difficult nodes through the principle of homophily [76] and utilizes curriculum learning to adjust the size of the matching window for trajectories. This approach allows GEOM to systematically structure the condensation procedure to progress from simpler to more challenging trajectories, thereby incorporating more diverse supervision signals from the original graph.
@@
Kernel ridge regression. Beyond aligning gradients or trajectories during model training, kernel ridge regression is employed for precise solutions in bi-level optimization objectives. For instance, GC-SNTK [61] discards the GNN and employs the kernel ridge regression, incorporating graph neural tangent kernels [77] as the relay model. This modification not only facilitates the first-order optimization of the condensed graph but also eliminates the requirement for multiple initializations of relay GNN, resulting in a more stable condensation procedure. KiDD [29] employs kernel ridge regression for graph classification tasks and further improves the method by removing the non-linear activation in graph neural tangent kernels. This approach simplifies the matrix multiplications in the iterative updates of condensed graphs and leads to substantial improvements in model accuracy compared to the gradient matching method.
4.1.2 Augmentation-Based Methods
@@
Optimization-driven methods enhance the effectiveness of GC by emphasizing the optimization procedure. In contrast, other studies improve the performance by augmenting different modules in the GC framework.
Structure augmentation. MSGC [48] and GCSR [60] improve the GC from graph structure construction. Unlike many GC approaches which synthesize a single dense adjacency matrix, MSGC instead pre-defines multiple small-scale sparse graphs. These pre-defined structures allow each condensed node to encompass distinct neighborhoods. The variation in structure enables GNNs to capture broader information, leading to diverse node embeddings and enhanced performance. Additionally, GCSR explicitly models the inter-class correlations within the original graph structure and leverages them to reconstruct an interpretable condensed graph structure through the self-expressive property [78]. This approach transfers structural correlations from the original graph to the condensed graph, thereby facilitating better information preservation and enhancing the graph structure.
Label augmentation. PUMA [65] and CGC [66] emphasize the significance of label quantity in GC and improve the condensation quality by incorporating pseudo labels from the original graph. The inclusion of additional supervision signals enhances class representation and ensures more accurate information preservation in the condensed graph.
Feature augmentation. GroC [47] designs a perturbation module in the process of generating condensed graphs. This module maximizes the distance between the condensed graph and the original graph by selectively perturbing the condensed node features that are insufficiently informative. By utilizing the adversarial training technique, these perturbations are capable of exploring the neighborhood areas within the parameter space of the condensed graph and significantly enhancing the gradient matching process. TinyGraph [57] aims to further compact the condensed graph by reducing the dimension of node features. It employs a graph attention network to compress the original graph features, with the compressed node features replacing the original ones as the condensation target during the condensation process. By simultaneously optimizing the graph attention network and the condensed graph with a unified GC objective, TinyGraph effectively reduces the size of condensed graphs while preserving the downstream GNN performances.
@@ 4.2 Generalized Graph Condensation
The essence of generalization lies in preserving crucial task-related information while mitigating bias injection into the condensed graph. Nevertheless, some GNN relay models, employed for encoding both structure and feature information, inevitably result in a loss of structural information and create an entanglement between the condensed graph and the relay model. These issues significantly affect the generalization ability to unseen models and tasks. To explore the cause, GCEM [64] conducts an investigation across various GNN architectures, including different spatial and spectral GNNs. They find that the performance gap increases when the relay model and downstream model have different architectures. This is due to that distinct spectral preferences of GNNs have varying impacts on the significance of different graph eigenvectors [2, 79], consequently leading to alterations in node features. For instance, if a GNN functions primarily as a low-pass filter, it will prioritize low-frequency information in node features, overlooking the full spectrum. The loss of high-frequency information will injure the high-pass filter GNNs’ performance.
@@
To address this issue, some generalized GC methods are developed from the spectral perspective. For example, GCEM [64] eschews the conventional relay GNN and directly generates the eigenbasis for the condensed graph, thereby eliminating the spectrum bias inherent in relay GNNs. To address the inconsistencies in eigenvector quantity and dimension, GCEM aligns the condensed eigenvectors with the most important eigenvectors of the original graph and matches the class features in the subspace defined by these eigenvectors. Apart from leveraging the advanced relay model, SGDD [30] employs the Laplacian energy distribution (LED) to analyze the structural properties of the graph, quantifying the spectral shift between condensed and original graphs. A substantial LED shift suggests that crucial structural information from the original graph is lost, diminishing the generalization capabilities of the condensed graph. Therefore, SGDD leverages the LED as a structural shift metric and designs an additional regularization to enhance the structure of the condensed graph. This regularization aligns the LED between the original graph and the condensed graph, facilitating the broadcast of structural information. By preserving crucial structural information, SGDD effectively reduces the performance discrepancies observed across different GNN architectures and various tasks, e.g. node classification, anomaly detection and link prediction.
Besides these spectral methods, OpenGC [62] emphasizes the temporal generalization capacity of condensed graphs, particularly in dynamic open-world scenarios [70]. In this context, new nodes and classes are continuously integrated into the existing graph structure, resulting in a distribution shift compared to the original graph being condensed. To adapt the condensed graph to these dynamic distribution shifts and manage graph data over an extended period, OpenGC employs temporal data augmentation to simulate evolving graph patterns and create multiple structure-aware environments. Furthermore, it incorporates invariance learning regularization in GC to preserve the invariant patterns across different temporal environments in the condensed graph. By this means, OpenGC significantly enhances the adaptability of downstream GNNs by training them on the temporally generalized condensed graph, eliminating the need to laboriously design specific generalization modules for each GNN.
4.3 Efficient Graph Condensation
The GC process requires significant computational resources. To comprehend the root causes of heavy computation, we break down the GC process into three fundamental stages, including efficient graph encoding, efficient optimization, and efficient graph generation, and examine acceleration methods for each component.
4.3.1 Efficient Graph Encoding
In GC, the graph encoding procedure is executed by the relay model, and it requires intensive computation primarily due to the neighbor explosion problem in the large original graph. Generally, relay models used in GC methods adopt either the message-passing framework or the graph neural tangent kernel framework. All these frameworks stack multiple propagation and transformation operations to aggregate and update information according to the graph structure. However, as multiple propagation processes are executed, the number of aggregated neighboring nodes increases exponentially, leading to significant computational demands.
To mitigate the expensive computation resulting from propagation, some GC methods borrow the idea from SGC [16] and remove the non-linear transformation layers between aggregation layers in the relay model. For example, GCond [25] and PUMA [65] utilize SGC to encode the original graph. KiDD [29] simplifies the graph neural tangent kernel by removing non-linear activation at certain layers. By this means, the propagation of these methods can be pre-computed in the pre-processing procedure and only needs to be executed once. Instead of performing propagation during each training epoch, the online time complexity of GC is significantly reduced than utilizing other relay models.
4.3.2 Efficient Optimization
In GC, the optimization procedure requires numerous iterations to update the relay model before optimizing the condensed graph, leading to considerable computation.
Gradient matching. To mitigate this problem, DosCond [52] proposes the one-step gradient matching strategy which discards the iterated relay model update and only performs gradient matching for one single step. The theoretical analysis posits that minimizing the one-step matching loss serves a dual purpose: it not only directs the learning process of condensed graphs towards lower losses on original graphs but also delineates the trajectory for updating the condensed data. Moreover, one-step matching eliminates the requirement of hyper-parameter tuning, such as the number of iterations of bi-level optimization, further simplifying the condensation process. Besides simplifying the relay model updates in the inner loop optimization, EXGC [54] aims to expedite the convergence of condensed graph generation in the outer loop. It begins by framing the bi-level optimization within an Expectation Maximization framework, then employs the Mean-Field variational approximation to update a subset of node features at once while maintaining complementary features fixed. Furthermore, EXGC observes that the condensation procedure follows a long-tail distribution, where a minority of condensed node features retain the majority of performance capabilities. However, this distribution leads to significant variability in the convergence speeds across different node sets. To address this, EXGC introduces post-hoc explanation methods to pinpoint the key subgraphs that hold most of the essential information for further convergence speedup.
Kernel ridge regression. To avoid the bi-level optimization in GC, another research line involves kernel ridge regression-based methods [29, 61, 62]. These methods reformulate the classification task as a kernel ridge regression task and incorporate the graph neural tangent kernel within kernel ridge regression to encode the original graph. Kernel ridge regression provides an efficient closed-form solution, dispensing with the necessity for iterative computations in updating the relay model and resulting in faster convergence speeds.
Distribution matching. Besides the gradient matching and kernel ridge regression, distribution matching changes the optimization principle towards approximating the original graph in the feature space, attracting significant attention in efficient GC studies. This approach leads to a shift in the role of the relay model within GC. It transitions from aligning task performances to transforming both original and condensed graphs into a unified space for distribution alignment. Consequently, distribution matching obviates the need to compute the gradients and trajectories w.r.t. parameters of the relay model in each training iteration, leading to a substantial reduction in computational complexity. Specifically, GCDM [63], CaT [27] and PUMA [65] compute node distributions for both original and condensed graphs in the same feature space and employ the maximum mean discrepancy [80] between each pair of class distributions as the optimization objective. GCEM [64] decomposes the graph and matches the class distributions within sub-spaces formed by eigenvectors. To enhance the fidelity of the condensation procedure, DisCo [67] and SimGC [68] introduce the pre-trained GNN in distribution matching and further align the node variations for performance enhancement.
Clustering. Instead of updating the condensed graph via gradients, CGC [66] simplifies the distribution matching objective by recasting it as a class partition problem, which can be effectively resolved using any clustering method. By integrating this strategy with a predefined graph structure, CGC facilitates a training-free solution for generating condensed graphs. This innovation significantly accelerates the condensation process, enhancing the efficiency and scalability of GC.
Computation tree preservation. Mirage [69] proposes a training-free solution for graph-level datasets. The underlying concept is that message-passing GNNs decompose the input graph into a multiset of computation trees according to the neighborhood aggregation. These computation trees capture the fundamental connectivity information in the graph and explicitly model the message-passing flows for each node. By preserving the most frequently occurring computation trees for each class, Mirage identifies the most crucial graph structures, thereby achieving a relay model-agnostic condensation method.
4.3.3 Efficient Graph Generation
Condensed graph generation targets two components: the feature matrix and the adjacency matrix. Modeling these matrices typically requires two distinct optimization processes. For example, GCond [25] models the condensed structure as a function of the condensed nodes, requiring the optimization of an additional adjacency matrix generator. To simplify the optimization procedure, MSGC [48] and CGC [66] leverage the pre-defined sparse graphs to eliminate the condensed graph optimization. GCSR [60] incorporates the self-expressive property to construct the graph, enabling a closed-form solution for the calculation of the adjacency matrix. Furthermore, several studies [25, 28, 63, 27, 65] discard the adjacency matrix directly and instead use an identity matrix to represent condensed structure. These structure-free GC methods implicitly encode the original graph’s topological structure into discriminative node features via relay GNNs and these enriched embeddings are informative to train different downstream GNNs.
@@ 4.4 Fair Graph Condensation
GC can effectively capture the essential task-specific characteristics of the original graph by exclusively optimizing for task performance. However, this task-centric optimization approach may introduce fairness concerns, primarily attributed to the indiscriminate transmission of information into the condensed graph. Especially in scenarios with intensive compression rates, GC tends to amplify biases as it prioritizes the preservation of predominant and representative information w.r.t the tasks. Consequently, GNNs trained on condensed graphs exhibit more severe fairness issues than those trained on original graphs [49, 50] and challenge the applicability of GC in high-stake applications [81, 49].
@@
To mitigate bias in condensed graphs, FGD [49] initially investigates the relationship between the feature spaces of the original graph and condensed graph, and constructs an estimator for sensitive features. Following this, the variance of the estimated sensitive group membership is employed as the bias measurement and integrated as the regularization to facilitate the unbiased condensed graph. Moreover, GCARe [50] directly regularizes the condensation process with adversarial training. It additionally incorporates a discriminator into the relay model, which is designed to predict the sensitive groups of node representations. The relay model operates as a generator, generating fair data to deceive the discriminator. Consequently, the condensation process is formulated as a min-max optimization problem, striking a balance between data generation and debiasing efforts.
While the two aforementioned methods address group fairness from distinct perspectives, their focus has primarily been on node classification tasks, leaving fairness concerns in other tasks largely unexplored. Furthermore, an empirical study by FGD has revealed that datasets also experience more pronounced issues with individual fairness when GNNs are trained on condensed graphs [49]. This underscores the need to broaden the scope of fairness research to include additional tasks and alternative fairness metrics, such as individual fairness [82] and rank fairness [83].
4.5 Robust Graph Condensation
GC typically assumes the availability of the clean original graph and emulates its distribution indiscriminately by the condensed graph. However, this idealized assumption contradicts the complexities of data processing in practical scenarios. Real-world graphs are often marred by various forms of noise—ranging from data collection errors [84] and outdated information [85] to inherent variability in dynamic environments [86, 87]. Additionally, recent research [88] has highlighted vulnerabilities of GC to backdoor attacks, where malicious attackers can subtly manipulate the condensed graph while maintaining its apparent quality. These unexpected noises or deliberate attacks all significantly impact the representation of the condensed graph and carry over to GNNs, compromising the model prediction accuracy and the resilience of GC in real-world applications.
Despite its significant relevance, robust GC has not been extensively studied, with only a limited number of studies addressing this vital field. RobGC [58] emerges as a pioneer by specifically targeting structural noises and incorporates a denoising procedure into the graph condensation process. It employs the condensed graph as a denoising signal to provide an overview for the structure and node features of the original graph, facilitating the correction of inherent noises. By alternately executing graph condensation and denoising procedures, the quality of both the original and condensed graphs is mutually improved, ensuring the condensed graph retains only the essential information from the original graph. Moreover, the unified optimization in RobGC enables the denoising of unseen test graphs using the noise-free condensed graph, leading to more accurate inference results.
As a preliminary exploration, RobGC effectively demonstrates the feasibility and crucial need for developing robust GC methods. However, its application is currently limited to addressing structural noises. Expanding the research scope to include challenges such as node feature noise and label noise is imperative for enhancing the utility of GC across a broader spectrum of practical applications.
TABLE III: Summary of structure modeling and sparsification methods in existing graph condensation research, with representative literature listed.
Methods Principles Literature
Structure modeling Generative model Node similarity GCond [25], SGDD [30], DisCo [67]
Parameterization
Node self-expression &
Inter-class structure correlation
GCSR [60]
Low rank decomposition KiDD [29]
Spectral decomposition GCEM [64]
Pre-defined structure Node similarity MSGC [48], CGC [66]
Structure-free method GCond [25], SFGC [28]
Structure sparsification Threshold - GCond [25]
Regularization - SGDD [30]
Gumbel softmax - DosCond [52], KiDD [29]